OpenStudy (anonymous):

INTEGRATE ABSOLUTE VALUE- 2x-x^2 from -4 t0 3

OpenStudy (anonymous):

this was my approach, x(2-x) =0 x =2 but then I ten to realise that as the numbers get to be >2, the fucntion is negative, same when they <2.....or am i missing something?

OpenStudy (anonymous):

integrated: x^2 - x^3/3 evaluate at 3 3^2 - 3^3/3 = 0 evaluate at -4 -4^2 + 64/3 = 16/3

ganeshie8 (ganeshie8):

i think you're trying to integrate this : \(\large \int\limits_{-4}^{3}|2x-x ^{2}|dx\) ?

OpenStudy (anonymous):

yes, I forgot the bars, thank you

ganeshie8 (ganeshie8):

x(2-x) =0 x = 0, x = 2 x >2 : negative x < 2 : positive x < 0 : negative

ganeshie8 (ganeshie8):

so, -4->0 : |2x-x^2| = x^2-2x 0->2 : |2x-x^2| = 2x - x^2 2 -> 3 : |2x-x^2| = x^2-2x

OpenStudy (anonymous):

wow, hey, sorry I was waayy off! i'll keep my mouth shut and learn from the masters!

ganeshie8 (ganeshie8):

you need to split it into 3 integrals

OpenStudy (anonymous):

ok, let me try that. because when its less than 2, its not at all positive

ganeshie8 (ganeshie8):

you're right it is positive only between 2 and 0 x < 0 is negative .... thats the reason, we need 3 integrals

OpenStudy (anonymous):

yes, i got 40. I went wolfram first, but then trying to figure out reverse wise was becoming difficult. Thank you, but then explain to me why we broke it into three.

ganeshie8 (ganeshie8):

good question :)

ganeshie8 (ganeshie8):

we broke it cuz we dont know how to take integral of absolute value

ganeshie8 (ganeshie8):